Calculate the amount of heat (in kJ) required to heat 255 g of water from 25.2 0C to 90.5 0C. (Cs = 4.18 J/g 0C for water)
Answer #1

The amount of heat required to heat m g of water from t1 0C to t2 0C; (q) = m*c*DT

m = mass of water = 255 g

c = specific heat of water = 4.18 j/g.c

DT = (90.5-25.2) 0C

The amount of heat (in J) required to heat 255 g of water from 25.2 0C to 90.5 0C, q = 255*4.18*(90.5-25.2) = 69603.27 J

The amount of heat (in KJ) required to heat 255 g of water from 25.2 0C to 90.5 0C, (q) = 255*4.18*(90.5-25.2) = 69.60327 KJ